# DriveSuprim: Towards Precise Trajectory Selection for End-to-End Planning

### v2

@@ 3.4 Self-distillation with Soft-labeling
During inference, the teacher model is utilized to output results.
4 Experiments
In this section, we first introduce our implementation details. Next, we show the superior performance of DriveSuprim on NAVSIM v1 and NAVSIM v2. Results of ablation studies are listed to validate the effectiveness of the proposed modules. Finally, we produce some visualization results to intuitively show the advantages of our method.
4.1 Implementation Details
Dataset and Metrics
We conduct experiments mainly on the NAVSIM dauner2024navsim benchmark, which is a driving dataset containing challenging driving scenarios.
There are two different evaluation metrics on NAVSIM, leading to NAVSIM v1 and NAVSIM v2. The evaluation metric of NAVSIM v1 is the PDM Score (PDMS). Each predicted trajectory is sent to a simulator to collect different rule-based subscores, which are aggregated to get the final PDMS:
$\mathrm{PDMS}=\left(\prod_{m\in S_{P}}{\mathrm{score}_{m}}\right)\times\left(\frac{\sum_{w\in S_{A}}{\mathrm{weight}_{w}\times\mathrm{score}_{w}}}{\sum_{w\in S_{A}}{\mathrm{weight}_{w}}}\right),$ (12)
where $S_{P}$ and $S_{A}$ denote the penalty subscore set and the weighted average subscore set. In NAVSIM v1, $S_{P}$ comprises two subscores, including no collisions (NC) and drivable area compliance (DAC), and $S_{A}$ comprises ego progress (EP), time-to-collision (TTC), and comfort (C). NAVSIM v2 introduces 4 more subscores, including driving direction compliance (DDC) and traffic light compliance (TLC) in $S_{P}$, and lane keeping (LK) and extended comfort (EC) in $S_{A}$. The comfort subscore is also revised to become the history comfort (HC). Aggregating all these subscores leads to the $\mathrm{EPDMS}$ metric in NAVSIM v2.
Model Details
We conduct our methods on three different backbones as the image encoder $\mathrm{Enc}_{i}$, including ResNet34 he2016deep , VoVNet lee2019an , and ViT-Large dosovitskiy2021an . The input images are resized to a resolution of $2048\times 512$ for ResNet34 and VoVNet, and $1024\times 256$ for Vit-Large. The ViT-Large backbone is pre-trained through Depth Anything yang2024depth . A 2-layer MLP is leveraged as $\mathrm{Enc}_{t}$ to encode each trajectory in the vocabulary. The trajectory decoder $\mathrm{TransDec}$ and refinement decoder $\mathrm{RefineDec}$ are both 3-layer Transformer Decoders with 256 hidden dimensions. The prediction head $\mathrm{head}^{(m)}$ predicts the normalized $l_{2}$ distance to the human ground-truth trajectory, and each subscore in NAVSIM except for EC. We adopt a 2-layer MLP for the distance prediction, and a linear layer for predicting other metrics. The number of trajectories in the vocabulary and filtered trajectories of the coarse filtering stage are set to 8192 and 256. We adopt a 3-camera FOV setting, with images from $l_{0}$, $f$, and $r_{0}$ cameras as the input. The rotation angle boundary $\Theta$ is set to $\pi/6$ in the augmentation pipeline. The threshold $\delta_{m}$ in soft-labeling is 0.15.
@@ 4.1 Implementation Details
Transfuser chitta2022transfuser ResNet34 97.7 92.8 79.2 92.8 100 84.0
UniAD hu2023planning ResNet34 97.8 91.9 78.8 92.9 100 83.4
PARA-Drive weng2024para ResNet34 97.9 92.4 79.3 93.0 99.8 84.0
VADv2 chen2024vadv2 ResNet34 97.9 91.7 77.6 92.9 100 83.0
LAW li2024enhancing ResNet34 96.4 95.4 81.7 88.7 99.9 84.6
DRAMA yuan2024drama ResNet34 98.0 93.1 80.1 94.8 100 85.5
Hydra-MDP li2024hydra ResNet34 98.3 96.0 78.7 94.6 100 86.5
DiffusionDrive diffusiondrive ResNet34 98.2 96.2 82.2 94.7 100 88.1
@@
Hydra-MDP li2024hydra V2-99 98.4 97.8 86.5 93.9 100 90.3
DriveSuprim V2-99 98.0 98.2 90.0 94.2 100 92.1 (+1.8)
Hydra-MDP li2024hydra ViT-L 98.4 97.7 85.0 94.5 100 89.9
@@
Training Details
We train our model on 8 NVIDIA A100. We adopt the Adam optimizer for model training, the batch size on a GPU is 8, and the learning rate is set to $7.5\times 10^{-5}$. More training details about the training pipeline are shown in appendix A.
4.2 Main Results
Result on NAVSIM v1
We evaluate the performance of DriveSuprim on the NAVSIM v1 benchmark. As shown in Tab 2, our method reaches 89.9% PDMS with the ResNet34 backbone, surpassing DiffusionDrive by 1.8%. Moreover, DriveSuprim with the stronger Vit-Large backbone can reach 93.5% PDMS, which is close to the human agent.
Result on NAVSIM v2
Table 3 shows that our model can also reach the SOTA result on the more challenging NAVSIM v2 benchmark. On the EPDMS metric, DriveSuprim surpasses previous SOTA methods by 1.7%, 0.9%, and 1.5%, respectively.
@@ 4.2 Main Results
Transfuser chitta2022transfuser ResNet34 96.9 89.9 97.8 99.7 87.1 95.4 92.7 98.3 87.2 76.7
HydraMDP++ li2024hydramdp_pp ResNet34 97.2 97.5 99.4 99.6 83.1 96.5 94.4 98.2 70.9 81.4
@@
Multi- stage Aug Data Self- distill NC $\uparrow$ DAC $\uparrow$ DDC $\uparrow$ TL $\uparrow$ EP $\uparrow$ TTC $\uparrow$ LK $\uparrow$ HC $\uparrow$ EC $\uparrow$ EPDMS $\uparrow$
98.4 98.6 99.6 99.8 90.5 97.8 97.0 98.3 78.6 87.1
98.7 98.5 99.5 99.8 87.5 98.3 96.0 98.3 71.1 85.6 (-1.5)
98.2 98.3 99.5 99.8 90.3 97.6 96.4 98.3 77.9 86.4 (-0.7)
97.5 98.3 99.5 99.8 90.7 96.8 96.4 98.3 75.4 85.6 (-1.5)
@@
Single-stage 98.5 98.5 99.5 99.7 87.4 97.9 95.8 98.2 75.7 85.6
+ 6 layer decoder 98.3 98.3 99.4 99.8 87.5 97.8 95.2 98.3 75.9 85.3 (-0.3)
+ Layer-wise scoring 98.7 98.4 99.4 99.9 87.2 98.1 95.5 98.2 75.6 85.6 (+0.3)
+ Trajectory filtering 98.6 98.2 99.5 99.9 89.3 98.0 97.0 98.3 76.2 86.4 (+0.8)
Table 6: Ablation on refinement setting. “Stage Layer” is the layer number of $\mathrm{RefineDec}$, and “Top-K” denotes the number of trajectories selected by the coarse filtering stage.
Stage Layer Top-K EPDMS $\uparrow$
1 256 86.6
3 256 87.1
6 256 86.6
3 64 86.5
3 512 86.9
3 1024 86.4
Table 7: Results with different soft label thresholds. $\delta_{m}$ denotes the soft label threshold in Equ 9.
$\delta_{m}$ EPDMS $\uparrow$
0.00 86.8
0.15 87.1
0.30 86.8
0.70 86.5
1.00 86.6
4.3 Ablation Studies
@@
We conduct ablation studies on the modules and approaches we propose, the result is shown in Tab 4. Adopting multi-stage refinement leads to a remarkable 1.5% performance gain on PDMS, and the augmented data improves the EPDMS by 0.7%. Applying the self-distillation framework is also beneficial to the training, with a 1.5% EPDMS improvement.
@@
To thoroughly validate the effectiveness of our coarse-to-fine trajectory selection paradigm, we devise a series of model evolutions from single-stage selection to coarse-to-fine selection, to eliminate the effect of parameter numbers in the model decoder. The results are listed in Tab 5. Firstly, we expand the number of trajectory decoder layers from 3 to 6, aligned with our coarse-to-fine setting, then we adopt the layer-wise scoring to the decoder. Nevertheless, these two revisions don’t lead to performance improvement. Only after we decrease the number of trajectories in the refinement stage does the performance increase from 85.6% to 86.4%.
@@
We further conduct comprehensive ablation studies on the refinement approach and soft-labeling. the results are shown in Tab 7 and Tab 7. Utilizing a 3-layer refinement Transformer decoder and filtering 256 trajectories for the refinement stage leads to the best result. For the teacher soft label, choosing 0.15 as the threshold leads to the best EPDM score.
Comparison of different FOVs
Our model performance on different FOV settings is shown in appendix C.1. Utilizing images from 3 camera sensors leads to the best result.
4.4 Visualization and Analysis
@@
Tab. 8 demonstrates the superior performance of DriveSuprim in turning scenarios. We divide the test dataset into three subsets based on the turning angle of the ground-truth trajectories: $\mathrm{NAVTEST}_{\mathrm{left}}$ (left turns exceeding 30 degrees), $\mathrm{NAVTEST}_{\mathrm{forward}}$ (near-straight trajectories), and $\mathrm{NAVTEST}_{\mathrm{right}}$ (right turns exceeding 30 degrees). The performance improvements are more pronounced in turning scenarios than in near-straight ones, highlighting DriveSuprim ’s enhanced ability to handle turning maneuvers.
Table 8: EPDMS on three dataset splits. $\mathrm{NAVTEST}_{\mathrm{left}}$, $\mathrm{NAVTEST}_{\mathrm{forward}}$, and $\mathrm{NAVTEST}_{\mathrm{right}}$ involve left-turning scenarios, near-forward scenarios, and right-turning scenarios.
Method $\mathrm{NAVTEST}_{\mathrm{left}}$ $\mathrm{NAVTEST}_{\mathrm{forward}}$ $\mathrm{NAVTEST}_{\mathrm{right}}$
Hydra-MDP++ li2024hydramdp_pp 68.7 87.2 77.7
DriveSuprim 71.6 (+2.9) 88.1 (+0.9) 79.7 (+2.0)
@@
Fig. 4 illustrates the trajectory frequency distribution in both the original dataset and the dataset enhanced with our rotation-based augmentation method. Using the ground-truth PDM score, we count the trajectory that either scores above 0.99 or ranks among the top-three-highest scores in each scenario. We then normalize the frequency by setting the most frequent trajectory to 1. Results show that in the original dataset, trajectories are predominantly concentrated in the forward or near-forward direction, whereas in the augmented dataset, trajectories across all directions appear with similar frequency.
Visulization results
The visualization results are shown in appendix B. This qualitative result shows that DriveSuprim not only performs well in complex challenging scenarios, but also handles sharp turn scenarios perfectly.
5 Conclusion
In this work, we present DriveSuprim, a novel framework for end-to-end autonomous driving that addresses three key challenges of selection-based methods. By introducing a coarse-to-fine selection strategy and an integrated training pipeline including rotation-based data augmentation and soft-label self-distillation, DriveSuprim significantly enhances the model’s ability to distinguish hard negatives to make more nuanced decisions to precisely select trajectories, and performs well in scenarios involving sharp turns. Extensive experiments on the NAVSIM benchmark demonstrate that our approach outperforms prior methods by a substantial margin.
Appendix A Training Details and Inference Pipeline
@@
During inference, we compute a final score for each trajectory by linearly combining the predicted scores across different metrics. We select the filtered and final trajectory based on the final score. Specifically, for the imitation learning metric and the multiplier metrics introduced in NAVSIM dauner2024navsim , we apply a coefficient to the logarithm of the predicted score. For the weighted average metric in NAVSIM, the coefficient is applied to the predicted score, followed by an additional logarithmic process of the sum. The overall inference score can be formulated as:
@@
Appendix B Visualization Results
Fig 5 shows qualitative visualization results comparing our method with the selection-based approach Hydra-MDP++ li2024hydramdp_pp . The subimages in the first row are the results of Hydra-MDP++, and the second row is our result.
The subimages in the first column show a challenging driving scenario where the ego vehicle is preparing to overtake a leading vehicle just before a crossroad. DriveSuprim successfully executes the overtaking maneuver, whereas the classic selection-based method incorrectly does a left turn, potentially leading to a collision. The subimages in the second column demonstrate DriveSuprim’s ability to distinguish between trajectories that appear right at first glance, i.e., the “hard negatives”. Although both models generate similar trajectories, the trajectory produced by Hydra-MDP++ deviates from the road centerline at the endpoint, while our method generates a near-perfect trajectory that closely matches the human expert’s ground truth. The third and fourth columns highlight the superior performance of our model in sharp-turning scenarios. DriveSuprim consistently generates smooth and accurate turning trajectories even under large turning angles.

### v3

@@ Methods
During inference, the teacher model is utilized to output the planning trajectories.
@@
Transfuser (2022) ResNet34 97.7 92.8 79.2 92.8 100 84.0
UniAD (2023) ResNet34 97.8 91.9 78.8 92.9 100 83.4
VADv2 (2024) ResNet34 97.9 91.7 77.6 92.9 100 83.0
LAW (2024b) ResNet34 96.4 95.4 81.7 88.7 99.9 84.6
DRAMA (2024) ResNet34 98.0 93.1 80.1 94.8 100 85.5
Hydra-MDP (2024c) ResNet34 98.3 96.0 78.7 94.6 100 86.5
DiffusionDrive (2025) ResNet34 98.2 96.2 82.2 94.7 100 88.1
@@
Hydra-MDP (2024c) V2-99 98.4 97.8 86.5 93.9 100 90.3
DriveSuprim V2-99 98.0 98.2 90.0 94.2 100 92.1 (+1.9)
Hydra-MDP (2024c) ViT-L 98.4 97.7 85.0 94.5 100 89.9
@@
Transfuser (2022) ResNet34 96.9 89.9 97.8 99.7 87.1 95.4 92.7 98.3 87.2 76.7
HydraMDP++ (2024a) ResNet34 97.2 97.5 99.4 99.6 83.1 96.5 94.4 98.2 70.9 81.4
@@
Method DS $\uparrow$ SR $\uparrow$ Eff. $\uparrow$ Comf. $\uparrow$
DriveAdapter (2023) 64.22 33.08 70.22 16.01
Hydra-NeXt (2025b) 73.86 50.00 197.76 20.68
Orion (2025) 77.74 54.62 151.48 17.38
AutoVLA (2025) 78.84 57.73 146.93 39.33
DriveSuprim 83.02 60.00 238.78 20.89
Table 4: Evaluation on Bench2Drive.
Experiments
In this section, we first introduce our implementation details. Next, we show the superior performance of DriveSuprim on NAVSIM and Bench2Drive, and ablation studies on NAVSIM are listed to validate the effectiveness of the proposed modules. Finally, we produce some visualization results to intuitively show the advantages of our method.
Implementation Details
Dataset and Metrics
We conduct experiments mainly on the NAVSIM (Dauner et al. 2024) and Bench2Drive (Jia et al. 2024a) benchmark. NAVSIM includes two evaluation metrics, leading to NAVSIM v1 and NAVSIM v2. The evaluation metric of NAVSIM v1 is the PDM Score (PDMS). Each predicted trajectory is sent to a simulator to collect different rule-based metrics, which are weighted aggregated to get the final PDMS. NAVSIM v1 includes 5 metrics: no collisions (NC), drivable area compliance (DAC), ego progress (EP), time-to-collision (TTC), and comfort (C). NAVSIM v2 further introduces 4 extended metrics, including driving direction compliance (DDC), traffic light compliance (TLC), lane keeping (LK) and extended comfort (EC). Aggregating all these subscores leads to the EPDMS metric.
Bench2Drive is a closed-loop benchmark evaluating 220 short routes in CARLA v2. The evaluation metrics include Success Rate, Driving Score, Efficiency and Comfortness. More details about datasets and metrics are in Appendix A.
Model Details and Training Details
We conduct our methods on three different backbones as the image encoder $\mathrm{Enc}_{i}$, including ResNet34 (He et al. 2016), VoVNet (Lee et al. 2019), and ViT-Large (Dosovitskiy et al. 2021). A 2-layer MLP is leveraged as $\mathrm{Enc}_{t}$ to encode each trajectory in the vocabulary. The trajectory decoder $\mathrm{TransDec}$ and refinement decoder $\mathrm{RefineDec}$ are both 3-layer Transformer Decoders with 256 hidden dimensions. The MLP prediction head $\mathrm{head}^{(m)}$ predicts the normalized $l_{2}$ distance to the human ground-truth trajectory, and each subscore in NAVSIM except for EC. The number of trajectories in the vocabulary and filtered trajectories of the coarse filtering stage are set to 8192 and 256. We adopt a 3-camera FOV setting, with images from $l_{0}$, $f$, and $r_{0}$ cameras as the input. The rotation angle boundary $\Theta$ is set to $\pi/6$ in the augmentation pipeline. The threshold $\delta_{m}$ in soft-labeling is set to 0.15.
We train our model on 8 NVIDIA A100. We use Adam for model training, the batch size on a GPU is 8, and the learning rate is set to $7.5\times 10^{-5}$. More model details and training details are shown in Appendix B.
Main Results
Result on NAVSIM
Tab 2 shows the performance of DriveSuprim on the NAVSIM benchmark. our method reaches 89.9% PDMS with the ResNet34 backbone, surpassing DiffusionDrive by 1.8%. Moreover, DriveSuprim with a stronger Vit-Large backbone can reach 93.5% PDMS. Results in Table 3 show that our model can also reach the SOTA result on the more challenging NAVSIM v2 benchmark. On the EPDMS metric, DriveSuprim surpasses previous SOTA methods by 1.7%, 0.9%, and 1.5%, respectively. DriveSuprim reaches 27.2 FPS and 12.5 FPS with ResNet34 and ViT-Large backbone, achieving real-time planning capability.
Result on Bench2Drive
Evaluation on Bench2Drive shows the close loop planning capability of our method. Based on the CARLA-Garage dataset (Jaeger, Chitta, and Geiger 2023; Sima et al. 2024) and the TF++ framework (Jaeger, Chitta, and Geiger 2023), our approach adopts the two-stage trajectory prediction for longitudinal control. Tab 4 shows that DriveSuprim achieves 83.02 and 60.00 scores on Driving Score and Success Rate, with 238.78 Efficiency and 20.89 Comfortness, outperforming the previous SOTA significantly.
Multi-
stage
Aug
Data
Self-
distill
97.2 97.5 99.4 99.6 83.1 96.5 94.4 98.2 70.9 81.4
97.5 96.4 99.1 99.6 87.0 96.4 95.3 98.2 75.0 82.4
96.9 96.9 99.4 99.6 87.9 96.0 95.5 98.3 76.5 82.7
97.5 96.5 99.4 99.6 88.4 96.6 95.5 98.3 77.0 83.1
@@
EPDMS $\uparrow$
Single-stage 81.4
+ 6 layer decoder 81.7
+ Layer-wise scoring 81.9
+ Trajectory filtering 82.4
@@
Ablation Studies
@@
We conduct ablation studies on the modules and approaches we propose, and the result is shown in Tab 5. Compared to baseline, adopting multi-stage refinement leads to 1.0% performance gain on PDMS, and the improvements further gained by augmented data and the self-distillation framework are 0.3% and 0.4%.
@@
We evolve our model from conventional single-stage selection to coarse-to-fine selection, to eliminate the effect of parameter number increase and validate the effectiveness of the coarse-to-fine trajectory selection paradigm. Tab 6 reveals that increasing the number of decoder layers from 3 to 6 brings only 0.3% improvement in EPDMS, while introducing layer-wise scoring and trajectory filtering leads to a more substantial 0.7% gain, proving that the performance boost comes from the coarse-to-fine mechanism rather than model size increase.
Figure 4: Visualization results across various challenging scenarios. In each example, the green trajectory represents the ground truth from the human expert, the red trajectory is generated by Hydra-MDP++, and the blue trajectory is produced by DriveSuprim.
@@
We further conduct comprehensive ablation studies on the refinement approach and soft-labeling. Utilizing a 3-layer refinement Transformer decoder and filtering 256 trajectories for the refinement stage leads to the best result. For the teacher soft label, choosing 0.15 as the threshold leads to the best EPDM score. Results are shown in Appendix C.
Method $\mathrm{NAVTEST}_{\mathrm{l}}$ $\mathrm{NAVTEST}_{\mathrm{f}}$ $\mathrm{NAVTEST}_{\mathrm{r}}$
Hydra-MDP++ 68.7 87.2 77.7
DriveSuprim 71.6 (+2.9) 88.1 (+0.9) 79.7 (+2.0)
Table 7: EPDMS on three dataset splits. $\mathrm{NAVTEST}_{\mathrm{l}}$, $\mathrm{NAVTEST}_{\mathrm{f}}$, and $\mathrm{NAVTEST}_{\mathrm{r}}$ involve left-turning scenarios, near-forward scenarios, and right-turning scenarios.
Visualization and Analysis
Visualization Results
Fig 4 shows qualitative visualization results on NAVSIM comparing DriveSuprim with the selection-based approach Hydra-MDP++. The images in the first and second columns are the results of Hydra-MDP++ and DriveSuprim. The first row illustrates a challenging overtaking scenario near a crossroad, where DriveSuprim correctly overtakes while Hydra-MDP++ mistakenly turns left and risks collision. In the second row, DriveSuprim outperforms in a sharp turn with smooth and accurate trajectories. These results demonstrate DriveSuprim not only performs well in challenging scenarios by choosing precise trajectories, but also excels in handling sharp turns with high accuracy. More visualization results are shown in Appendix D.
@@
Tab. 7 demonstrates the superior performance of DriveSuprim in turning scenarios. We divide the test dataset into three subsets based on the turning angle of the ground-truth trajectories: $\mathrm{NAVTEST}_{\mathrm{l}}$ (left turns exceeding 30 degrees), $\mathrm{NAVTEST}_{\mathrm{f}}$ (near-forward trajectories), and $\mathrm{NAVTEST}_{\mathrm{r}}$ (right turns exceeding 30 degrees). Performance improvements of DriveSuprim are more pronounced in turning scenarios than in near-straight ones, highlighting enhanced ability to handle turning maneuvers.
@@
We illustrate the trajectory frequency distribution in Appendix D. Results show that in the original dataset, trajectories are predominantly concentrated in the forward or near-forward direction, while in the augmented dataset, trajectories across all directions appear with similar frequency.
Conclusion
We present DriveSuprim, a novel framework for end-to-end planning. By introducing coarse-to-fine selection and an integrated training pipeline with rotation-based data augmentation and soft-label self-distillation, DriveSuprim significantly enhances the model’s ability to distinguish hard negatives and precisely select trajectories, and performs well in scenarios involving sharp turns. Extensive experiments on the NAVSIM and Bench2Drive benchmark demonstrate that our approach outperforms prior methods by a substantial margin.
Appendix
A. More information about Datasets and Metrics
NAVSIM
NAVSIM (Dauner et al. 2024) is split into navtrain and navtest, containing 103k and 12k samples for model training and testing. There are two different evaluation metrics on NAVSIM, leading to NAVSIM v1 and NAVSIM v2. The evaluation metric of NAVSIM v1 is the PDM Score (PDMS). Each predicted trajectory is sent to a simulator to collect different rule-based subscores, which are multiplied or weighted aggregated to get the final PDMS:
$\mathrm{PDMS}=\left(\prod_{m\in S_{P}}{\mathrm{s}_{m}}\right)\times\left(\frac{\sum_{w\in S_{A}}{\mathrm{w}_{w}\times\mathrm{s}_{w}}}{\sum_{w\in S_{A}}{\mathrm{w}_{w}}}\right),$ (12)
where $S_{P}$ and $S_{A}$ denote the multiplier subscore set and the weighted average subscore set, $\mathrm{s}$ denotes the subscore, and $\mathrm{w}$ is the weight of each subscore. In NAVSIM v1, $S_{P}$ comprises two subscores: no collisions (NC) and drivable area compliance (DAC), and $S_{A}$ comprises ego progress (EP), time-to-collision (TTC), and comfort (C). NAVSIM v2 introduces 4 more subscores, including driving direction compliance (DDC) and traffic light compliance (TLC) in $S_{P}$, and lane keeping (LK) and extended comfort (EC) in $S_{A}$. The comfort subscore is also revised to become the history comfort (HC). Aggregating all the above subscores leads to the $\mathrm{EPDMS}$ metric.
Bench2Drive
Bench2Drive (Jia et al. 2024b) is a recently proposed closed-loop evaluation benchmark for end-to-end autonomous driving, which includes 220 short routes spanning 44 interactive driving scenarios. Each route is around 150 meters. This short-distance route design allows for a comprehensive and accurate analysis of the planning model’s capability. The evaluation metrics include the Success Rate (SR), Driving Score (DS), Efficiency, and Comfortness. The Success Rate measures the proportion of routes executed correctly by the planner, the Driving Score considers the planner’s performance from route completion and rule compliance, the Efficiency checks whether the vehicle speed is too slow, and the Comfortness measures the smoothness by comparing the trajectory with human behaviors.
B. Supplementary Implementation Details
Model details
We conduct our methods on three different backbones as the image encoder $\mathrm{Enc}_{i}$, including ResNet34 (He et al. 2016), VoVNet (Lee et al. 2019), and ViT-Large (Dosovitskiy et al. 2021). The input images are resized to a resolution of $2048\times 512$ for ResNet34 and VoVNet, and $1024\times 256$ for Vit-Large. The ViT-Large backbone is pre-trained through Depth Anything (Yang et al. 2024).
Training and inference pipelines
@@ B. Supplementary Implementation Details
During inference, we compute a final score for each trajectory by linearly combining the predicted scores across different metrics, and the filtered and final trajectories are selected based on the final score. Specifically, for the imitation learning metric and the multiplier metrics introduced in NAVSIM (Dauner et al. 2024), we apply a coefficient to the logarithm of the predicted score. For the weighted average metric in NAVSIM, the coefficient is applied to the predicted score, followed by an additional logarithmic process of the sum. The overall inference score can be formulated as:
@@
C. Supplementary Experiment Result
Ablation on refinement and soft label settings
We further conduct comprehensive ablation studies on the refinement approach and soft-labeling. Tab 10 shows that utilizing a 3-layer refinement Transformer decoder and filtering 256 trajectories for the refinement stage leads to the best result. Tab 11 and Tab 12 show ablations on soft label. For the teacher soft label, choosing 0.15 as the clipping threshold and adopting a 1-meter shift in soft imitation learning lead to the best EPDM score.
Stage
Layer
Top-K
EPDMS $\uparrow$
1 256 83.0
3 256 83.1
6 256 82.6
3 64 81.8
3 512 82.9
3 1024 83.0
Table 10: Ablation on refinement setting. “Stage Layer” is the layer number of the Refinement Decoder, and “Top-K” denotes the number of trajectories selected by the coarse filtering stage.
$\delta_{m}$ EPDMS $\uparrow$
0.00 82.6
0.15 83.1
0.30 83.0
0.70 82.9
1.00 82.8
Table 11: Results with different soft label thresholds. $\delta_{m}$ denotes the soft label threshold in the Equation 9.
$d_{i}$ EPDMS $\uparrow$
0.0 82.6
0.5 83.0
1.0 83.1
2.0 82.7
Table 12: Results with different soft imitation trajectory shift. $d_{i}$ denotes the soft imitation trajectory shift in $L_{\mathrm{imi-soft}}$ in the Equation 10.
Model performance with different FOV settings
We show the model performance with different FOV settings in Tab 15. The FOV is revealed by the number of cameras. As shown in Fig 5, five cameras are involved in these settings: $f$ (front camera), $l_{0}$ (front-left camera), $l_{1}$ (left camera), $r_{0}$ (front-right camera), and $r_{1}$ (right camera). The 1-camera setting uses $f_{0}$ as input, the 3-camera setting uses $l_{0}$, $f_{0}$ and $r_{0}$ as input, while the 5-camera setting uses $l_{1}$, $l_{0}$, $f_{0}$, $r_{0}$ and $r_{1}$ as input. Experimental results show that the 5-camera setting leads to the best performance.
Superiority of self-distillation
DriveSuprim utilizes a self-distillation framework to mitigate the optimization challenges posed by hard binary ground-truth labels. We compare our self-distillation framework with other smoothing techniques to validate its superiority. We adopt temperature-scaled cross-entropy and vallina label smoothing, respectively. The temperature is set to 2.0, and the label smoothing coefficient is set to 0.1. As shown in Tab 13, our proposed self-distillation framework achieves the best performance on EPDMS.
Parameter number and inference time
We list the number of parameters and inference time of DriveSuprim in Tab 14. Compared to Hydra-MDP++, our model introduces only a minor increase of approximately 7M parameters. Moreover, the FPS drop due to the parameter increase is acceptable. DriveSuprim maintains the real-time planning capability.
D. Visualization
More visualization result
Fig 6 shows qualitative visualization results comparing our method with the selection-based approach Hydra-MDP++ (Li et al. 2024a). In each example, the image with a grey border shows the BEV visualization, the image with a red border is the result of Hydra-MDP++, and the image with a blue border is our result.
The first three columns on the left show challenging driving scenarios with complex vehicle interaction. In the first row, the ego vehicle attempts to overtake a leading vehicle just before a crossroad. DriveSuprim successfully completes the overtaking maneuver, whereas the classic selection-based method incorrectly does a left turn, potentially leading to a collision. The second and third rows further highlight DriveSuprim’s ability to distinguish between trajectories that appear right at first glance, i.e., the “hard negatives”. Although both models generate similar trajectories, the trajectories produced by Hydra-MDP++ deviate from the road centerline at the endpoint, whereas our method generates near-perfect trajectories that closely align with the human expert’s ground truths. The last three columns highlight the superior performance of our model in sharp-turning scenarios. DriveSuprim consistently produces smooth and accurate turning trajectories even under large turning angles.